### v2

# Meta-Learning Objectives for Preference Optimization

Abstract
Evaluating preference optimization (PO) algorithms on LLM alignment is a challenging task that presents prohibitive costs, noise, and several variables like model size and hyper-parameters. In this work, we show that it is possible to gain insights on the efficacy of PO algorithm on much simpler benchmarks. We design a diagnostic suite of MuJoCo tasks and datasets, which we use to systematically evaluate PO algorithms, establishing a more controlled and cheaper benchmark. We then propose a novel family of PO algorithms based on mirror descent, which we call Mirror Preference Optimization (MPO). Through evolutionary strategies, we search this class to discover algorithms specialized to specific properties of preference datasets, such as mixed-quality or noisy data. We demonstrate that our discovered PO algorithms outperform all known algorithms in the targeted MuJoCo settings. Finally, based on the insights gained from our MuJoCo experiments, we design a novel PO algorithm that significantly outperforms existing baselines in an LLM alignment task.
1 Introduction
Learning from human preferences (Christiano et al., 2017) is a paradigm which enables the alignment of machine learning systems to relative human preferences, without requiring access to absolute rewards. While the framework was developed for robotic and games applications with experiments on MuJoCo simulations and Atari (Akrour et al., 2012; Biyik & Sadigh, 2018; Ibarz et al., 2018), this paradigm has been successfully applied to Large Language Models (Team et al., 2023; Achiam et al., 2023). In particular, fine-tuning pre-trained LLMs with human preferences has become a popular strategy to adapt them to specific tasks and to improve their safety and helpfulness.
Within this framework, Reinforcement Learning from Human Feedback (RLHF) is one of the most popular methods. It consists in learning a reward function using a preference dataset and then optimizing the estimated reward using Reinforcement Learning methods such as Proximal Policy Optimization (PPO)  (Schulman et al., 2017). However, the training pipeline of RLHF is quite complex, which is why implicit approaches such as Direct Preference Optimisation (DPO)  (Schulman et al., 2017) have gained traction thanks to their simplicity. These methods do not learn a reward model but estimate it implicitly using the policy of the agent. Many follow ups to DPO have been proposed  (Yuan et al., 2023; Zhao et al., 2023; Azar et al., 2024; Xu et al., 2024a; Hong et al., 2024; Park et al., 2024; Meng et al., 2024), but comparing their performance in LLM alignment is a complex task that incurs high costs, noise, and the inherent difficulty in judging a response better than another.
In this work, we provide a comprehensive analysis of PO algorithms, examining their behavior on automatically generated preference datasets. We return to the roots of RLHF by performing this analysis in MuJoCo environments and datasets, where the underlying ground-truth reward structure is well defined and offers a clear performance metric to compare agents. In particular, we design a task where a pre-trained agent has to adhere to a new stylistic constraint, emulating the typical conditions of LLM fine-tuning. Our findings indicate that many PO algorithms present distinct failure modes when applied to specific mixed-quality or noisy datasets.
Moreover, we introduce a framework for finding PO algorithms. Specifically, we define a class of PO algorithms based on mirror descent (Nemirovski & Yudin, 1983), which generalizes DPO and ORPO for particular choices of the mirror map. We then show that this class can be easily parametrized and searched using evolutionary strategies (ES), optimizing for the final performance of the trained policy, as measured by the ground truth reward.
For each setting we consider, we discover an algorithm that significantly outperforms all baselines. Analyzing the discovered algorithms, we find that the main difference between them and the baselines is that they keep optimizing the policy of the agent well after the probability of generating the chosen trajectory has surpassed the probability of generating the rejected one. We use this insight to design a new PO function, Constrained Linear Policy Optimization (CoLPO) that demonstrates promising results in an LLM alingment task. We summarize our contributions below.
  1. 1.
    We perform a systematic evaluation of eight of the most popular PO algorithms on automatically generated preference datasets with varying levels of data quality, noise levels and initial policy. We see that most existing algorithms struggle when dealing with noise and mixed-quality date.
  2. 2.
    We introduce a novel family of offline PO algorithms using mirror descent, named Mirror Preference Optimization (MPO), which can be easily parameterized and explored via ES.
  3. 3.
    For both noisy and mixed-quality settings, we find and describe a PO algorithm within our framework that largely outperforms all the considered baselines in our MuJoCo benchmark.
  4. 4.
    We demonstrate that takeaways from our analysis on the MuJoCo setting, as well as the characteristics of the discovered PO algorithms, can be successfully transferred onto LLM tasks. In particular, we show that our CoLPO algorithm significantly improves upon the ORPO, CPO, and SLiC-HF baselines.
Overall, we show that it is possible to gain relevant insights for the design of LLM alignment algorithms through a systematic analysis in simple, computationally inexpensive environments such as MuJoCo.
2 Preliminaries
A policy $\pi\in(\Delta(\mathcal{A}))^{\mathcal{S}}$, where $\Delta(\mathcal{A})$ is the probability simplex over $\mathcal{A}$, represents the behavior of an agent on an MDP, whereby at state $s\in\mathcal{S}$ the agents takes actions according to the probability distribution $\pi(\cdot\mid s)$. Let $\tau=\{(s_{t},a_{t})\}_{t=0}^{T-1}$ denote a trajectory of length $T$ and, with a slight overload of notation, let $\pi(\tau)=\prod_{t=0}^{T-1}\pi(a_{t}|s_{t})$ and $r(\tau)=\sum_{t=0}^{T-1}r(s_{t},a_{t})$. Lastly, let $\pi(\cdot|\tau)$ be a distribution over $(\Delta(\mathcal{A}))^{T}$ defined as $\pi(\cdot|s_{0})\times\dots\times\pi(\cdot|s_{N-1})$.
Our objective is to find a policy $\pi^{\star}$ that maximizes the expected cumulative reward of an episode, that is
$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\tau\sim(\mu,\pi,P)}r(\tau)$ (1)
$:=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mu,a_{t},s_{t+1}}\sum_{t=0}^{T-1}r(s_{t},a_{t}),$
where $a_{t}\sim\pi(\cdot|s_{t})$ and $s_{t+1}\sim P(\cdot|s_{t},a_{t})$.
Let $\mathcal{D}=\{(s_{0}^{i},\tau_{w}^{i},\tau_{l}^{i})_{i=1}^{N}\}$ be a preference dataset, where each tuple $(s_{0},\tau_{w},\tau_{l})$ consists of a starting state $s_{0}$ and two trajectories with starting state $s_{0}$. Each pair of trajectories is ranked by a judge, who determines a chosen trajectory $\tau_{w}$ (“win”) and a rejected trajectory $\tau_{l}$ (“lose”), based on the cumulative rewards $r(\tau_{w})$ and $r(\tau_{l})$. Most settings assume the judge ranks trajectories according to the Bradley-Terry model (Bradley & Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
$\mathbb{P}(\tau_{w}\succ\tau_{l})$ $=\frac{\exp(r(\tau_{w})/\eta)}{\exp(r(\tau_{w})/\eta)+\exp(r(\tau_{l})/\eta)}$ (2)
$=\sigma((r(\tau_{w})-r(\tau_{l}))/\eta),$
where $\sigma$ is the sigmoid function and $\eta$ is a temperature parameter. In this work, we consider an offline training setting, where the agent aim to solve the optimization problem in (1) but only has access to the the dataset $\mathcal{D}$ and cannot collect further data. We also assume the agent does not have access to either the transition probability $P$, the reward function $r$, or the MDP $\mathcal{M}$.
2.1 Alignment to preference feedback
There are several algorithms in the literature to optimize the objective in (1) using a preference dataset $\mathcal{D}$. We describe supervised fine-tuning (SFT), DPO and ORPO, as they are among the most popular and as many methods can be seen as a variation of one of these algorithms.
SFT
SFT is an initial alignment phase, where the policy $\pi_{0}$ is trained to imitate high-quality demonstration data. In particular, the starting policy $\pi_{0}$ is updated to minimize the cross-entropy loss $\ell(\pi,(s_{0},\tau_{w},\tau_{l}))=-\log(\pi(\tau_{w}))$. We call reference policy $\pi_{\text{ref}}$ the policy obtained at the end of this procedure.
DPO
Direct Preference Optimization (DPO) consists in solving a maximum likelihood estimation problem and a policy optimization problem in a single step. The maximum likelihood estimation problem is the one to find an estimate of the true reward function that governs how the preferences are expressed, that is
$\widehat{r}\in\mathop{\mathrm{argmax}}_{r_{\theta}}g\circ\sigma((r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l}))/\eta),$ (3)
for a parametrized reward class $\{r_{\theta}:\theta\in\Theta\}$, for an increasing function $g:\mathbb{R}\leftrightarrow\mathbb{R}$, and for $\eta\geq 0$. Here, typically, $g$ is set to be the logarithm function. The policy optimization problem is the one to maximize the expected reward, that is
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}$ (4)
$\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}\widehat{r}(s_{t},a)-\beta D_{\mathtt{KL}}(\pi(\cdot|\tau),\pi_{\text{ref}}(\cdot|\tau))\right],$
where $D_{\mathtt{KL}}$ represents the $\mathtt{KL}$-divergence and is introduced to prevent the policy from moving too far away from the dataset distribution.
DPO merges these two problems by using the agent itself to implicitly represent the reward model. It consists in optimizing the objective
$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$ (5)
$\left[\log\sigma\left(\beta\left(\log\frac{\pi(\tau_{w})}{\pi_{\mathrm{ref}}(\tau_{w})}-\log\frac{\pi(\tau_{l})}{\pi_{\mathrm{ref}}(\tau_{l})}\right)\right)\right],$
which is obtained by plugging the theoretical solution of (4) in the maximum likelihood problem in (3), with $g=\log$. Refer to Appendix C for details. Thanks to its simplicity, DPO has been widely adopted to fine-tune LLMs (Yuan et al., 2024; Jiang et al., 2024).
A known issue of DPO is that it pushes probability mass away from the preference dataset and to unseen responses (Xu et al., 2024b), which can cause the final policy to deviate significantly from the reference policy, even when the reference policy aligns well with human preferences. To mitigate this risk, DPO is usually applied for a few epochs.
ORPO
ORPO further simplifies the training pipeline and addresses the distribution shift issue present in DPO. It merges the SFT and DPO steps into one, optimizing the unified objective
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\bigg{[}\underbrace{\log\pi(\tau_{w})}_{\text{SFT}}$ (6)
$+\lambda\underbrace{\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)}_{\text{preference optimization}}\bigg{]}$
where $\operatorname{odds}_{\pi}(\tau)=\pi(\tau)/(1-\pi(\tau))$. ORPO gets rid of the need for a reference model by adding an SFT term to the preference optimization objective function, and uses this term to prevent the optimized policy from moving too far away from the dataset distribution. Additionally, the SFT term prevents pushing probability mass away from the preference dataset, addressing the distribution shift issue present in DPO.
Research on preference optimization has been very active and many methods have been proposed. We present a summary of the most popular algorithms in Table 1 and a brief discussion in Appendix A.
Beyond these implicit algorithms, there are several other methods that explicitly solve the maximum likelihood problem in (3) and use the learned reward model to optimize the objective in (4) with an RL algorithm. Overall, RLHF is a superior approach and the industry standard, but is more computationally expensive and complex to implement. For a detailed discussion and comparison between DPO-like methods and PPO, refer to Appendix F.
Method Objective
RRHF (Yuan et al., 2023) $\lambda\log\pi_{\theta}(\tau_{w}|x)-\max\left(0,-\frac{1}{|\tau_{w}|}\log\pi_{\theta}(\tau_{w}|x)+\frac{1}{|\tau_{l}|}\log\pi_{\theta}(\tau_{l}|x)\right)$
SLiC-HF (Zhao et al., 2023) $\lambda\log\pi_{\theta}(\tau_{w}|x)-\max(0,\delta-\log\pi_{\theta}(\tau_{w}|x)+\log\pi_{\theta}(\tau_{l}|x))$
DPO (Rafailov et al., 2024) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\beta\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}\right)$
IPO (Azar et al., 2024) $-\left(\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}-\frac{1}{2\tau}\right)^{2}$
CPO (Xu et al., 2024a) $\log\pi_{\theta}(\tau_{w}|x)+\log\sigma(\beta\log\pi_{\theta}(\tau_{w}|x)-\beta\log\pi_{\theta}(\tau_{l}|x))$
ORPO (Hong et al., 2024) $\log\pi(\tau_{w})+\lambda\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)$
R-DPO (Park et al., 2024) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\beta\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}+(\alpha|\tau_{w}|-\alpha|\tau_{l}|)\right)$
SimPO (Meng et al., 2024) $\log\sigma\left(\frac{\beta}{|\tau_{w}|}\log\pi_{\theta}(\tau_{w}|x)-\frac{\beta}{|\tau_{l}|}\log\pi_{\theta}(\tau_{l}|x)-\gamma\right)$
Table 1: A summary of popular PO algorithms. The objective in each row is to be maximized and $(\tau_{w},\tau_{l})\sim\mathcal{D}$.
2.2 Mirror Maps
We review the concept of mirror map, which will be needed when describing our methodology. For a convex set $\mathcal{X}\subseteq\mathbb{R}^{|\mathcal{A}|}$, a mirror map $h:\mathcal{X}\rightarrow\mathbb{R}$ is defined as a strictly convex, continuously differentiable and essentially smooth function [^1] A function $h$ is essentially smooth if $\lim_{x\rightarrow\partial\mathcal{X}}\lVert\nabla h(x)\rVert_{2}=+\infty$, where $\partial\mathcal{X}$ denotes the boundary of $\mathcal{X}$. function that satisfies $\nabla h(\mathcal{X})=\mathbb{R}^{|\mathcal{A}|}$. Essentially, a mirror map is a function whose gradient allows bijective mapping between the primal space $\mathcal{X}$ and the dual space $\mathbb{R}^{|\mathcal{A}|}$. The specific class of mirror maps that we are going to use is the $\omega$-potential mirror map class, to which most mirror maps considered in the literature belong.
Definition 2.1 ($\omega$-potential mirror map (Krichene et al., 2015)).
For $u\in(-\infty,+\infty]$, $\omega\leq 0$, an $\omega$-potential is defined as an increasing $C^{1}$-diffeomorphism $\phi:(-\infty,u)\rightarrow(\omega,+\infty)$ such that
[…]
For any $\omega$-potential $\phi$, we define the associated mirror map $h_{\phi}$ as
$h_{\phi}(\pi(\cdot|s))=\sum\nolimits_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}\phi^{-1}(x)dx.$
When $\phi(x)=e^{x-1}$ we recover the negative entropy mirror map, while we recover the $\ell_{2}$-norm when $\phi(x)=2x$ (refer to Appendix D). Mirror maps in this class are simple to implement in practice, where $\mathcal{A}$ is often large, as they can be parametrized by a scalar function instead of a multi-dimentional one. Additionally, the same $\omega$-potential $\phi$ can be used to generate mirror maps for different action spaces, allowing the insights obtained for one action space to easily generalize to others. An $\omega$-potential mirror map $h_{\phi}$ induces a Bregman divergence (Bregman, 1967), which is defined as
$\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))$ $:=h_{\phi}(\pi(\cdot|s))-h_{\phi}(\pi^{\prime}(\cdot|s))$
$-\langle\nabla h_{\phi}(\pi^{\prime}(\cdot|s)),\pi(\cdot|s)-\pi^{\prime}(\cdot|s)\rangle,$
where $\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\geq 0$ for all $x,y\in\mathcal{Y}$. When $\phi(x)=e^{x-1}$, $\mathcal{D}_{h_{\phi}}$ is equivalent to the KL-divergence, while we recover the Euclidean distance when $\phi(x)=2x$ (refer to Appendix D). When the Bregman divergence is employed as a regularization term in optimization problems, tuning the mirror map allows us to control the geometry of the updates of the parameters to be optimized, determining when to take large or small updates based on the current value of the parameters.
2.3 Evolution Strategies
OpenAI-ES  (Salimans et al., 2017) is a popular method to be able to optimize non-differentiable functions and it has been widely used to meta-learn objectives  (Lu et al., 2022; Jackson et al., 2024), as it obtains an unbiased estimate of the gradient (unlike second order gradient methods). The gradient $\nabla_{\zeta}F(\zeta)$ is estimated using:
$\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I_{d})}\left[\frac{\epsilon}{2\sigma}(\widehat{F}(\zeta+\sigma\epsilon)-\widehat{F}(\zeta-\sigma\epsilon))\right],$
where $\mathcal{N}(0,I_{d})$ is the multivariate normal distribution, $d$ is the number of parameters, $\widehat{F}$ is an estimate of $F$, and $\sigma>0$ is a hyperparameter regulating the variance of the perturbations.
3 Mirror Preference Optimization
We introduce Mirror Preference Optimization (MPO), a new framework for preference optimization that generalizes DPO and ORPO. We start by replacing the $\mathtt{KL}$-divergence penalty term in the objective in (4) with a Bregman divergence and aim to solve the problem
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}$ (7)
$\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}r(s_{t},a)-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\right],$
where $D_{h}$ is the Bregman divergence induced by a mirror map $h$. This new objective allows us to enforce different types of regularization, which, as we show later in the paper, can be tailored to account for specific properties of the preference dataset. Following the same intuition used to obtain the DPO objective, we have the following result.
Theorem 3.1.
Let $h_{\phi}$ be a 00-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
$r(\tau)=\phi^{-1}(\pi^{\star}(\tau))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau))+c(s_{0}),$ (8)
where $c(s_{0})$ is a normalization constant that depends only on $s_{0}$.
We provide a proof for Theorem 3.1 in Appendix C. By plugging (11) in the maximum likelihood problem in (3), we obtain the objective:
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\mathcal{D}}\big{[}g\circ\sigma\big{(}\beta(\phi^{-1}(\pi(\tau_{w}))$ (9)
$-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))\big{)}\big{]},$
where $\mathbb{E}_{\mathcal{D}}$ is equivalent to $\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$.
Two-step MPO
We can use the objective in (9) to define a class of two-step PO algorithms, which consist of a preliminary SFT phase to obtain the reference policy $\pi_{\mathrm{ref}}$ and a PO phase which optimizes (9). When $\phi=e^{x}$, the (9) is equivalent to (5) and we recover the DPO algorithm.
One-step MPO
By adding an SFT term to (9) and by setting the $\pi_{\mathrm{ref}}$ be the uniform distribution, we obtain a class of one-step PO algorithms. These algorithms consists in a single phase, where we optimize the objective
$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\psi(\pi(\tau_{w}))$ (10)
$+\lambda g\circ\sigma\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\big{]},$
where $\psi$ is an $\omega$-potential. Here, the terms $-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))$ and $\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))$ have canceled out due to $\pi_{\mathrm{ref}}$ being uniform. We note that setting $\pi_{\mathrm{ref}}$ to be the uniform distribution is equivalent to replacing the Bregman divergence penalty in (7) with the mirror map $h(\pi(\cdot|\tau))$, which enforces a form of entropy regularization. When $\psi(x)=\log(x)$ and $\phi^{-1}(x)=\log(x/(1-x))$, (10) recovers the ORPO objective in (6).
The objectives in (9) and (10) allow us to implement a wide variety of preference optimization algorithms, while benefiting from a theoretical justification. In the following, we will show that it is possible to parametrize and optimize $g$, $\psi$, and $\phi^{-1}$ to obtain new algorithms that outperform current baselines.
3.1 Meta Learning PO objectives
To search the space of PO algorithms we have defined, we employ a neural network parametrization for $g$, $\psi$, and $\phi^{-1}$, which we optimize using evolutionary strategies (Salimans et al., 2017).
Similarly to (Alfano et al., 2024), we parameterize $g$, $\psi$ and $\phi^{-1}$ as a one layer neural network with 126 hidden units and non-negative kernels, where the activation functions are equally split among:
$x,\;(x)_{+}^{2},\;x^{3},\;(x)_{+}^{1/2},\;(x)_{+}^{1/3},\;\log((x)_{+}),$
$e^{x},\;\tanh(x),\;\log(\mathop{\mathrm{clip}}(x)/(1-\mathop{\mathrm{clip}}(x))),$
where $(x)_{+}=\max(x,0)$ and $\mathop{\mathrm{clip}}(x)=\max(\min(x,1),0)$. The non-negative kernels and the increasing activation functions guarantee the monotonicity of $g$, $\psi$, and $\phi^{-1}$, while the several different activation functions facilitate expressing complex functions. To ensure that we are able to recover the DPO and ORPO objectives, we add $a\log(x)$, $b\log(x)$ and $c\log(x/(1-x))$ to the final outputs of $g$, $\psi$ and $\phi^{-1}$, respectively, where $a,b,c\geq 0$.
To search for the best $g$, $\psi$ and $\phi^{-1}$ within this class, we employ the OpenAI-ES strategy. Denote by $\zeta$ the parameters of $g$, $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$. Lastly, let $F(\zeta)$ be the expected cumulative reward of $\pi^{\zeta}$, i.e. $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$. We then use Adam (Kingma & Ba, 2015) to update the parameters $\zeta$ using the estimated gradient. In practice, to compute (12), we sample 128 values of $\epsilon$ to obtain 256 perturbed objective functions. We then train 256 agents with the perturbed objective functions on an a preference dataset. To measure the value of each agent, i.e. $\widehat{F}(\zeta^{\prime})$ for all perturbed $\zeta^{\prime}$, we sample 100 trajectories on the target environment for each agent, and take the average cumulative reward as estimate for the value of the agent. Refer to Appendix E for further discussion and details on the ES methodology.
In our experiments, we consider both the case where we fix $g=\log$ and learn $\psi$ and $\phi^{-1}$, and the case where we learn all three functions. We also perform the evolution on both the two-step and the one-step MPO classes.
Base Mixed Quality Noisy ($\varepsilon=0.1$) Noisy ($\varepsilon=0.2$) Noisy ($\varepsilon=0.3$)
RRHF 2789 $\pm$ 285 2245 $\pm$ 134 1730 $\pm$442 749 $\pm$ 498 330 $\pm$ 552
SLiC-HF 3255 $\pm$ 66 2478 $\pm$ 54 2329 $\pm$289 1964 $\pm$ 116 1135 $\pm$ 224
DPO 3528 $\pm$ 58 2766 $\pm$ 89 3082 $\pm$80 2530 $\pm$ 97 1519 $\pm$140
IPO 3618 $\pm$ 44 2937 $\pm$ 85 3162 $\pm$ 66 2392 $\pm$136 1133 $\pm$ 115
CPO 3450 $\pm$ 55 2322 $\pm$ 208 2967 $\pm$ 58 2427 $\pm$ 44 2000 $\pm$35
ORPO 3087 $\pm$ 322 2500 $\pm$ 71 2841 $\pm$ 50 2359 $\pm$ 43 1953 $\pm$ 37
R-DPO 2606 $\pm$ 65 2107 $\pm$ 40 2099 $\pm$ 50 1740 $\pm$ 36 1667 $\pm$ 24
SimPO 3683 $\pm$78 3117 $\pm$ 185 2314 $\pm$ 752 118 $\pm$ 715 -3828 $\pm$ 341
SFT 3287 $\pm$ 62 2344 $\pm$ 40 2733 $\pm$45 2345 $\pm$ 37 2049 $\pm$33
LPO 3774 $\pm$ 102 2841 $\pm$ 46 3617 $\pm$ 69 2705 $\pm$ 370 1569 $\pm$ 156
With $g=\log$
1S-MPO (mixed-quality) 3206 $\pm$ 330 3153 $\pm$ 274 1319 $\pm$ 714 -1625 $\pm$ 944 -3967 $\pm$ 382
1S-MPO (noisy, $\varepsilon=0.1$) 3789 $\pm$ 60 3210 $\pm$ 60 3813 $\pm$ 47 3280 $\pm$ 83 3279 $\pm$ 83
2S-MPO (mixed-quality) 3595 $\pm$ 57 2785 $\pm$ 78 3197 $\pm$ 58 2487 $\pm$ 110 1687 $\pm$ 58
2S-MPO (noisy, $\varepsilon=0.1$) 3551 $\pm$ 58 2784 $\pm$ 63 3190 $\pm$ 62 2552 $\pm$ 94 1569 $\pm$ 122
With parametrized $g$
1S-MPO (mixed-quality) 3560 $\pm$ 333 3627 $\pm$ 79 3371 $\pm$ 410 3230 $\pm$ 259 2681 $\pm$ 251
2S-MPO (mixed-quality) 3736 $\pm$ 51 3202 $\pm$ 64 3488 $\pm$ 73 2992 $\pm$ 87 2253 $\pm$ 125
1S-MPO (noisy, $\varepsilon=0.1$) 3861 $\pm$ 79 3075 $\pm$ 87 3724 $\pm$ 59 2845 $\pm$ 365 1771 $\pm$ 107
2S-MPO (noisy, $\varepsilon=0.1$) 3701 $\pm$ 52 3178 $\pm$ 59 3490 $\pm$ 95 2886 $\pm$ 127 2074 $\pm$ 136
1S-MPO (noisy, $\varepsilon=0.3$) 3931 $\pm$ 69 3244 $\pm$ 55 3834 $\pm$ 82 3735 $\pm$ 84 3417 $\pm$ 82
Table 2: Three Legged Ant (TLA) Adaptation. Performance of popular and discovered MPO algorithms on TLA, for various dataset settings. For each algorithm-dataset combination, we report the average value and standard error of 25 trained agents. For discovered MPO algorithms, we report performance of one-step MPO (1S-MPO) and two-step MPO (2S-MPO). For each discovered MPO algorithm, we specify on which setting it was discovered and report its performance across all settings (with fixed hyperparameters). We underline the highest (or two highest if their confidence interval overlaps) average performance among the human-designed algorithm and report in bold the overall highest, for each setting.
4 MuJoCo Experiments
Our first set of experiments is carried out on continuous RL tasks in MuJoCo. In particular, we show the performance of all the algorithms presented in Table 1 across several settings and we compare it with the performance of our discovered objectives. To maximize computational efficiency, all our MuJoCo experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries. We report our hyper-parameters in Appendix I.
4.1 Tasks
To reproduce the typical conditions of LLM fine-tuning, which involve a pre-trained model, we consider a setting where the task is to adapt a pre-trained agent to meet the original objective while adhering to an additional stylistic constraint. Specifically, in the Ant environment, we start from an agent that has been pre-trained on the standard Ant goal of moving forward and enforce the objective of avoiding the use of one of its legs. This is accomplished by introducing the Three-legged-ant (TLA) environment, a modified version of Ant where utilizing the fourth leg results in significant penalties.
The offline preference optimization task is defined as follows. We train one agent (the original agent) in the original Ant environment, achieving a reward of 6000, and another (the target agent) in the TLA environment, which achieves a reward of 3900. For comparison, the original agent achieves a reward of 1700 in the TLA environment. We then generate a preference dataset of 1280 rows, each with two trajectories of length 1000 starting from the same state. Each trajectory is generated by either the original or the target agent, depending on the current setting. A Bradley-Terry judge ranks each pair of trajectories and declares a winner, based on their true cumulative reward.
We also consider training a randomly initialized agent on the Hopper environment, using a preference dataset with 5120 pairs of trajectories. This setting addresses the case where a new behavior has to be learned completely from the preference dataset and no prior knowledge of the task is available. Due to space constraints, we report the results of the experiments for this task in Appendix H.
We consider three variations of the preference dataset, each meant to represent a common issue of real world data.
  • Base dataset: for each pair of trajectories, one is generated by the original agent and one by the target agent.
  • Noisy dataset: same as the base dataset but each label given by the judge is flipped with probability $\varepsilon$.
  • Mixed-quality dataset: each trajectory in the dataset is generated by an agent selected at random between the original and the target one. The resulting dataset will consist of, approximately, $25\%$ comparisons between two trajectories from the target agent, $50\%$ comparisons between a trajectory of the target agent and one of the original agent, and $25\%$ comparisons between two trajectories of the original agent.
Figure 1: SimPO
Figure 2: Discovered 1S-MPO (shuffled)
Figure 3: ORPO
Figure 4: Discovered 1S-MPO (noisy)
Figure 5: Plots of the absolute value of the gradient of SimPO, ORPO, and of two discovered objectives, i.e. 1S-MPO (shuffled) and 1S-MPO (noisy, $\epsilon=0.3$). The dots are sampled datapoints from the training distribution, with lighter points representing more recent datapoints during the course of training. In the bottom right and corner of each plot, SimPO’s and ORPO’s gradients go to zero, while the gradient of our discovered loss functions keep increasing. The effect of this on training is evident observing the datapoints: in ORPO and SimPO, the likelihood of the rejected points is not minimized, leading to decreased performance.
4.2 Results
We provide the results of our experiments for TLA in  Table 2, which reports the performance of several PO algorithm and of our discovered objectives. We performed a hyperparameter search for each algorithm-dataset combination and only report the performance of the best hyperparameters. All algorithms are run for 12 epochs over the preference dataset, with the exception of DPO, IPO and R-DPO, which are run for 2 epochs after 10 epochs of SFT.
First of all, we notice that none of the human-designed algorithms manages to recover the performance of the target agent and that most of them experience a large drop in performance in the mixed-quality and noisy settings.
Importance of SFT
The first group within Table 2 shows that SFT plays a key role in the performance across different settings. SimPO, which does not contain an SFT term nor an SFT step, is at the top of the leaderboard on the base and the mixed-quality setting but performs poorly on all noisy settings. IPO and DPO, which do not contain an SFT term but have an SFT step, are among the top performers on the base, mixed-quality and low noise settings. Their performance finally drops when the noise level reaches 0.3. Lastly, the algorithms that present an SFT term in their objectives, e.g. CPO, ORPO and, obviously, SFT, exhibit a subotpimal performance in the base and mixed-quality settings but are much more robust to noise than the other algorithms.
Keep optimizing
Furthermore, while RRHF and SliC-HF are very similar, SliC-HF allows the PO part of the objective to be clipped when $\pi(\tau_{w})>\pi(\tau_{l})+\delta$, rather than when $\pi(\tau_{w})>\pi(\tau_{l})$. This modification leads to a higher performance on all tasks, demonstrating that it is beneficial to keep optimizing the policy even if $\pi(\tau_{w})>\pi(\tau_{l})$. To further stress this point, we consider the objective
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\lambda\log\pi_{\theta}(\tau_{w}|x)-\log\pi_{\theta}(\tau_{l}|x)\big{]},$
which we call Linear Preference Optimization (LPO). LPO corresponds to SLiC-HF with $\delta=+\infty$ and obtains better results than most of the existing algorithms in Table 2, confirming that it is important to design objectives that do not flatten when $\pi(\tau_{w})>\pi(\tau_{l})$.
Discovered objectives
Table 2 also reports the performance of our discovered objectives, for both the case where we set the monotonic transformation $g$ to be the logarithmic function and where we parametrize and learn it. We learn a separate objective for each dataset setting and report the performance of each learned objective on all settings.
Differently from the human-designed algorithms, the discovered objectives recover the performance of the target agent in multiple instances and are more robust to the mixed-quality and noisy settings. We note that allowing the evolution procedure to learn $g$ leads to a better performance in all dataset settings. Additionally, we have that the objectives discovered within the two-step MPO class always have a lower performance than those within the one-step MPO class. This is probably due to the ability to modify the SFT term in the one-step MPO class, which is not present in the two-step class.
When exploring the one-step MPO class with $g=\log$, our discovery procedure always recovers a variation of CPO. That is, we obtain an objective that can be approximated as
$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\alpha\log(\pi(\tau_{w}))$
$+\lambda\log\sigma\left(\beta\log(\pi(\tau_{w}))-\beta\log(\pi(\tau_{l}))\right)\big{]},$
where the coefficients $\alpha$ and $\beta$ depend on the setting. In particular, we have a low value for $\alpha$ and a high value for $\beta$ in the noisy settings, while we observe the opposite in the mixed-quality setting. These results confirm the observations made on the hand-crafted objectives, whereby objectives with an SFT term are more robust to noise and objectives without are more robust to mixed-quality trajectories. When we search the two-step MPO class with $g=\log$, we recover DPO.
Figure 5 provides a visualization of the gradient of some of the objectives discovered when we parametrize and meta-learn $g$. In particular, we show the objectives discovered within the one-step MPO class on the mixed-quality and noisy ($\varepsilon=0.3$) settings. For comparison, we provide the same plots for the ORPO and SimPO objectives. The hand-crafted algorithms present a larger gradient when $\pi(\tau_{w})<\pi(\tau_{l})$ and a smaller one when $\pi(\tau_{w})>\pi(\tau_{l})$, that is, they induce large updates when the data-point contradicts the current behaviour of the agent, and small otherwise.
The objective discovered on the noisy dataset has the opposite behavior, meaning that it only optimizes the more robust-to-noise SFT term when $\pi(\tau_{w})<\pi(\tau_{l})$. As $\log(\pi(\tau_{w}))-\log(\pi(\tau_{l}))$ becomes larger, it shifts toward increasingly large updates thanks to the PO term. A similar pattern appears in the mixed-quality dataset, where the objective also increases its updates as the difference $\log(\pi(\tau_{w}))-\log(\pi(\tau_{l}))$ grows. The key distinction between these two losses is that the shuffled loss triggers high-gradient updates even when $\pi(\tau_{w})<\pi(\tau_{l})$. We highlight once more that both discovered objectives advocate to keep optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$.
For additional results and analysis for the Hopper task, we refer the reader to Appendix H.
WR $s_{\text{WR}}$ LCWR $s_{\text{LCWR}}$ avg length
SLiC-HF 16.77 1.32 13.18 0.28 3016
ORPO 24.47 1.52 19.25 0.29 1679
CPO 25.96 1.54 21.07 0.25 1623
CoLPO 31.94 1.65 24.01 0.21 2391
Table 3: Alpaca Eval LLM results. We report Win Rate (WR), standard error for Win Rate ($s_{\text{WR}}$), Length Controlled Win Rate (LCWR), standard error for LCWR ($s_{\text{LCWR}}$) and average length of responses. We trained gemma-7b on dpo-mix-7k for 3 epochs. We then compared our model to GPT-4 using Llama-3.3-70B-Instruct as judge. All algorithms use the same $\beta=0.5$ and the same learning rate of $55e-$ 5. In the case of CoLPO and SLiC-HF, we used the same cutoff rate at 5.
5 Experiments: LLM transfer
We show that the insights obtained in the MuJoCo environments can be transferred to the LLM alignment setting. In particular, we found that all our discovered loss functions kept optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, while all algorithms in Table 1 have a PO term whose gradient quickly reaches 0 when $\pi(\tau_{w})>\pi(\tau_{l})$.
Using this insight, we design a new PO objective, which we call Constrained Linear Preference Optimization (CoLPO) and define as
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\log\pi_{\theta}(\tau_{w}|x)$
$+\beta\mathbf{1}_{\{\log\pi_{\theta}(\tau_{w}|x)<\alpha\}}\min(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})),0)$
$+\beta\mathbf{1}_{\{\log\pi_{\theta}(\tau_{w}|x)>\alpha\}}\min(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})),\gamma)\big{]},$
where $\beta\geq 0$, $\gamma\geq 0$ and $\alpha\leq 0$ are hyperparameters. The CoLPO objective keeps optimizing the policy when $\pi(\tau_{w})>\pi(\tau_{l})$, but only if $\log\pi_{\theta}(\tau_{w}|x)$ is over a certain threshold. This constraint is placed to stop the optimizer from increasing the difference $\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))$ by descreasing both probabilities.
We evaluate CoLPO on an LLM fine-tuning task, where we fine-tune the gemma-7b (Team et al., 2024) model on the dpo-mix-7k [^2] https://huggingface.co/datasets/argilla/dpo-mix-7k dataset for 3 epochs. Being a real-world preference dataset, dpo-mix-7k naturally presents characteristics of the mixed-quality and the noisy settings. To tune the LLMs, we modify the Alignment Handbook library (Tunstall et al., ) to include the CoLPO objective. We evaluate the tuned LLM against GPT-4, using the AlpacaEval library  (Li et al., 2023). All models were trained using the default hyper-parameters, with $\beta=0.5$, $\gamma=5$ (for CoLPO and SLiC-HF), and $\alpha=-1$ (for CoLPO). For details on hyperparameters and training settings consult Appendix J.
Table 3 reports the results of CoLPO, as well as the results for other one-step algorithms, and shows that CoLPO largely outperforms existing algorithms, confirming the validity of the insights gained in the MuJoCo simulations.
Lastly, we report that the the log-probabilities for the final training batch of chosen and rejected responses are -1.17 and -11.38 respectively for CoLPO, compared to -0.61 and -2.98 for ORPO, respectively.
This results further confirms that CoLPO is effective in decreasing the probability of rejected responses and mantaining high the probability of chosen responses.
6 Conclusion
We have introduced a novel framework for Preference Optimization algorithms, as well as a methodology for the automatic discovery of PO algorithms using evolutionary strategies. Through a systematic evaluation across diverse settings in MuJoCo environments, we have demonstrated that the performance of our discovered objectives consistently exceeds the performance of existing methods, particularly in noisy and mixed-quality datasets where many baselines underperform.
Our analysis in MuJoCo also revealed a common shortcoming among current baselines: truncating the loss whenever $\pi(\tau_{w})>\pi(\tau_{l})$. To address this, we proposed a simple objective, CoLPO, that avoids such loss truncation while preventing the optimizer from decreasing the probability of chosen datapoints. We then tested this objective on an LLM fine-tuning task, achieving significant improvements over existing methods, thereby confirming the broader applicability of our approach.
Appendix A Popular PO Algorithms
These methods include RRHF, which uses length-normalized log-likelihood, and SLiC-HF, which uses direct log-likelihood and incorporates SFT.
The comparison also includes IPO, a theoretically-based approach that handles pairwise preferences differently than DPO. Another method, CPO, combines sequence likelihood as a reward with an SFT objective. Finally, R-DPO modifies the original DPO by adding regularization to prevent length exploitation.
Appendix B Related Work
Automatic Discovery of Preference Optimization Loss Functions
Several works in the literature have shown that it is possible to discover machine learning algorithms that outperform algorithms manually designed by researchers (Oh et al., 2020; Lu et al., 2022; Jackson et al., 2024; Alfano et al., 2024). An approach particularly relevant to our method is DiscoPOP by (Lu et al., 2024), which leverages an LLM to discover objective functions for LLM tuning. They consider a different space of objective functions from us, as they replace the log-sigmoid in (5) with a generic loss function, following the framework built by (Tang et al., 2024). Additionally, instead of searching over a space of parametrized functions, they ask the LLM to generate loss functions in code space. This distinction suggests that our approaches could be complementary, as the model discovered by DiscoPOP could be paired with our learned mirror map. Lastly, DiscoPOP optimizes its objective function directly on the final task, whereas we adopt a two-stage process—optimizing the loss function on a separate task (MuJoCo) and later transferring it to the LLM setting. This transferability underscores the broader applicability of our approach.
Generalisations of DPO
A generalization of DPO alternative to ours is $f$-DPO, developed by (Wang et al., 2023), which consists in replacing the $\mathtt{KL}$-divergence in (1) with an $f$-divergence and then apply the same heuristic as DPO to obtain the final objective function. We note that the $\mathtt{KL}$-divergence is the only $f$-divergence to be also a Bregman divergence, and vice-versa. They empirically demonstrate that different $f$-divergences lead to different balances between alignment performance and generation diversity, highlighting the trade-offs inherent to this class of algorithms. (Huang et al., 2024) further explore this class of PO algorithm and individuate an $f$-divergence for which $f$-DPO is robust to overoptimization.
Appendix C Proof of Theorem 3.1
We provide here a proof for our main result, i.e. Theorem 3.1. The proof to obtain the DPO objective in (5) follows by taking $\phi=e^{x}$.
Theorem C.1 (Theorem 3.1).
Let $h_{\phi}$ be a 00-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
$r(\tau)=\phi^{-1}(\pi^{\star}(\tau))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau))+c(s_{0}),$ (11)
for all trajectories $\tau$, where $c(s_{0})$ is a normalization constant that depends only on $s_{0}$.
Proof.
We use the KKT conditions to solve (7), i.e.
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\right]$
We use the stationarity condition to obtain the equation
$\nabla_{\pi(\tau)}$ $\Bigg{[}\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))$
$\quad-\lambda\left(\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\pi(\tau^{\prime})-1\right)+\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\alpha(\tau^{\prime})\pi(\tau^{\prime})\Bigg{]}$
$=r(\tau)-\beta\phi^{-1}(\pi(\tau))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau)=0,$
for all initial states $s_{0}\in\mathcal{S}$ and for all trajectories $\tau$ starting from $s_{0}$. Rearranging, we obtain that
$\pi(\tau)=\phi\big{(}r(\tau)+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau)\big{)}.$
Since $0\notin\mathop{\mathrm{dom}}\phi^{-1}$, due to the definition of a 00-potential, and $\pi_{\mathrm{ref}}(\tau)>0$, we have that $\pi(\tau)>0$ for all trajectories $\tau$. Invoking the complementary slackness condition, whereby $\alpha(\tau)\pi(\tau)=0$ for all trajectories $\tau$, we have that $\alpha(\tau)=0$ for all trajectories $\tau$. Therefore, we have that
$r(\tau)-\beta\phi^{-1}(\pi(\tau))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda=0$
The theorem statement is obtained by rearranging the last equation and denoting $c(s_{0})=\lambda$
Appendix D Further discussion of $\omega$-potentials
We show here two examples of Bregman divergence induced by an $\omega$-potential mirror map, that is when $\phi(x)=e^{x-1}$ and when $\phi(x)=x$. If $\phi(x)=e^{x-1}$, the associated mirror map is defined as
[…]
$=\sum_{a\in\mathcal{A}}\pi(a\mid s)\log(\pi(a\mid s))-\pi(a\mid s)+\pi(a\mid s)$
$=\sum_{a\in\mathcal{A}}\pi(a\mid s)\log(\pi(a\mid s)),$
which is the negative entropy. Plugging this expression in the definition of Bregman divergence we obtain
$\mathcal{D}_{h}(x,y)$ $=h(x)-h(y)-\langle\nabla h(y),x-y\rangle$
$=\sum_{a\in\mathcal{A}}x_{a}\log(x_{a})-y_{a}\log(y_{a})-(\log(y_{a})-y_{a})(x_{a}-y_{a})$
$=\sum_{a\in\mathcal{A}}x_{a}\log(x_{a}/y_{a}),$
which is the definition of the $\mathtt{KL}$-divergence. If $\phi(x)=2x$, the associated mirror map is defined as
[…]
$=\sum_{a\in\mathcal{A}}\pi(a\mid s)^{2},$
which is the $\ell_{2}$-norm. Plugging this expression in the definition of Bregman divergence we obtain
$\mathcal{D}_{h}(x,y)$ $=h(x)-h(y)-\langle\nabla h(y),x-y\rangle$
$=\sum_{a\in\mathcal{A}}x_{a}^{2}-y_{a}^{2}-(2y_{a})(x_{a}-y_{a})$
$=\sum_{a\in\mathcal{A}}(x_{a}-y_{a})^{2},$
which is the definition of the Euclidean distance.
Appendix E Further discussion on Evolution Strategies
Evolution Strategies (ES) represent a powerful, backpropagation-free method for optimizing complex functions, that has been particularly successful in the context of long-horizon, noisy, and bi-level optimization tasks such as RL and meta-RL. ES, and in particular the OpenAI-ES algorithm (Salimans et al., 2017), rely on perturbation-based sampling to estimate gradients without requiring backpropagation through the entire computational graph. This feature makes ES well-suited for tasks with long computational graphs, for instance algorithms with many updates, where, due to memory constraints, traditional gradient-based methods have to resort to gradient truncation, introducing bias (Werbos, 1990; Metz et al., 2022; Liu et al., 2022).
In our setting, we use ES to search for the best $\psi$ and $\phi^{-1}$ within the parametrized class introduced in Section 3.1, so that an agent trained using the objective in (10) achieves the highest value. Denote by $\zeta$ the parameters of $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$. Lastly, let $F(\zeta)$ be the expected cumulative reward (or value) of $\pi^{\zeta}$, i.e. $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$. At each iteration, we estimate the gradient $\nabla_{\zeta}F(\zeta)$ as
$\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I_{d})}\left[\frac{\epsilon}{2\sigma}(\widehat{F}(\zeta+\sigma\epsilon)-\widehat{F}(\zeta-\sigma\epsilon))\right],$ (12)
where $\mathcal{N}(0,I_{d})$ is the multivariate normal distribution, $d$ is the number of parameters, $\widehat{F}$ is an estimate of $F$, and $\sigma>0$ is a hyperparameter regulating the variance of the perturbations.
Appendix F Further discussion on Online vs Offline Methods
In the domain of RL and preference optimization, the choice between online and offline algorithms presents a critical trade-off, influencing computational efficiency, data requirements, and generalization capabilities. Online methods, such as PPO, iteratively collect and incorporate new data during training. These inherently support exploration of the environment, enabling the discovery of novel strategies or behaviors that are not captured in pre-existing datasets. However, they need feedback for each generated “trajectory” (or response, in the LLM case), which might be expensive to obtain. Online methods are also more complex and particularly sensitive to hyperparameters, often requiring meticulous tuning for stability and efficiency.
Offline algorithms, such as DPO and its variants, rely entirely on pre-collected datasets. These methods are designed for efficiency and simplicity: they don’t require any additional feedback from users and are therefore particularly effective in scenarios where feedback is delayed or unavailable. However, the reliance on static datasets means offline methods may struggle to generalize beyond the training data, particularly if the distribution shift between the training dataset and test time distribution is significant. Additionally, the performance of the algorithm is closely tied to the quality of the training dataset: noisy, biased, or corrupt datasets can severely degrade performance, as these methods cannot mitigate such issues through exploration or resampling.
In summary, RLHF (i.e., online) is considered the superior approach, particularly when substantial amounts of online labels are accessible. This makes it the industry standard (Xu et al., 2024b). While DPO has been theoretically equated to optimizing using PPO and a reward model trained on an offline dataset, recent empirical research (Tang et al., 2024) has challenged this notion. These studies have demonstrated that online methods, such as PPO, consistently outperform offline methods like DPO. This superiority is attributed to the benefits of on-policy sampling.
While DPO has occasionally outperformed PPO, it’s important to note that several studies (Xu et al., 2024b; Song et al., 2024) have consistently shown PPO’s overall superiority. DPO’s relative strength lies in its simpler training regime, which avoids the complexities associated with reward model inaccuracies. However, DPO’s performance is significantly limited by its sensitivity to distribution shift, especially when the offline preference data lacks diversity (Song et al., 2024). This limitation becomes particularly evident when querying the model with out-of-distribution data, a common challenge for methods relying solely on offline data. To mitigate this issue, DPO-iter (Xu et al., 2024b), which incorporates online data, has been proposed as a potential solution.
Appendix G Further discussion on Meta-Learning Algorithms
Meta-learning, or “learning to learn”, has been extensively employed to automate the design of algorithms that can either adapt rapidly with minimal data samples or generalize effectively to unseen data, tasks, or environments. The development of broadly applicable algorithms is particularly critical in the context of preference optimization for LLMs. Here, LLMs are fine-tuned on relatively small datasets of offline data but must generalize to a virtually infinite range of potential user queries. Prior work in meta-learning has demonstrated success in developing generalizable optimization algorithms and loss functions (Lu et al., 2022; Jackson et al., 2024; Lu et al., 2024; Goldie et al., 2024; Kirsch et al., 2020).
At its core, meta-learning is defined as a bilevel optimization problem with an inner and an outer loop. The inner loop consists in an iterative optimization algorithm that trains agents to solve a predetermined task given a set of meta-parameters. The outer loop consists in evaluating the agents trained in the inner loop and update the meta-parameters accordingly, following some optimization method like second order gradient descent (Finn et al., 2017). The evaluation of the agents is typically done on a held-out dataset in supervised learning or by sampling trajectories on the environment simulator in RL (Lu et al., 2022; Jackson et al., 2024). In our setting, the inner loop is the offline preference optimization algorithm, while the outer loop is the agent evaluation on the environment (online) and the update of the meta-parameters $\zeta$.
Base Mixed Quality Noisy $\epsilon=0.1$
DPO 1796 $\pm$ 78 458 $\pm$ 82 693 $\pm$ 131
IPO 2049 $\pm$ 13 1606 $\pm$ 91 739 $\pm$ 118
CPO 2078 $\pm$ 12 1078 $\pm$ 35 1813 $\pm$ 33
ORPO 2022 $\pm$ 15 1039 $\pm$ 20 1710 $\pm$ 33
SimPO 2027 $\pm$ 15 1460 $\pm$ 94 1794 $\pm$ 65
Table 4: Hopper. Perfomance of popular MPO algorithms on the Hopper setting. The agent is randomly initialised.
Appendix H MuJoCo Additional Results
Below, we report a simplified version of the discovered loss on the shuffled and noisy $\epsilon=0.3$ datasets:
$\mathcal{L}(\tau_{w},\tau_{l})=0.82\cdot\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$ (13)
$+1.4\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))$
$+0.22\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{2}$
$+0.2\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{3}$
$\mathcal{L}(\tau_{w},\tau_{l})=0.82\cdot\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$ (14)
$+\max((1.18\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{2},$
$0.1\cdot(1.18\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))))$
We consider a second set of simulations on MuJoCo, based on the Hopper environment. As for the previous sections, our experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries. We report our hyper-parameters in Appendix I.
H.1 Hopper Tasks
Differently from the TLA task, we consider a setting where the agent is randomly initialized and needs to learn a policy from scratch. On Hopper, we train an agent with an expected cumulative reward of 2100 (the expert agent) and an agent with an expected cumulative reward of 900 (the bad agent). We generate the preference datasets in the same way we do for TLA, with the exceptions that the number of rows is 5120 and the trajectories are generated by either the expert or the bad agent. We consider the same three variations of the preference datasets used in TLA, where the expert agent corresponds to the target agent, and the bad agent corresponds to the original agent. We include more data compared to the TLA setting as it takes more datapoints for the agent to learn from scratch rather than to adapt to a slightly different objective (like in the TLA case).
H.2 Further MuJoCo Analsys
In addition to the noisy dataset, we also considered a bad judge setting, where the judge would be more likely to swap the label of a pair of trajectories if their ground truth rewards were closer to each other. This is practically implemented as an increase in the temperature of the Bradley-Terry judge. However, we did not notice significantly different results compared to the simple noisy setting, therefore detailed results are not reported.
H.3 Results
Our Hopper results confirm our conclusions in the TLA setting. All algorithms but DPO come close to matching the performance of the expert agent (2100), with CPO being the best. We can see SimPO is the only algorithm that significantly outperforms the bad agent (performance of 900) in the Mixed Quality setting (the $\gamma$ for SimPO was set very high, $\gamma=10$, as a lower value significantly limited performance).
Appendix I MuJoCo Hyper-parameters
We give the hyper-parameters we use for training.
Parameter Value
Number of epochs 12
Minibatch size 2
Learning rate 1e-3
Max gradient norm 1.3
$\lambda$ 0.5
Table 5: Hyper-parameter settings for PO.
Parameter Hopper TLA
Population Size 256 256
Number of generations 128 256
Sigma init 0.03 0.03
Sigma Decay 0.999 0.999
Learning rate 0.02 0.02
Table 6: Hyper-parameter settings of OpenAI-ES.
Appendix J LLM Hyper-parameters
Parameter Value
Number of epochs 3
Gradient Accumulation Step 32
Batch Size 2
Total Batch Size 64
Learning rate 5e-5
LoRA Yes
LoRA Rank 128
LoRA Alpha 256
Lora Dropout 0.05
Max length 2048
Table 7: Hyper-parameter settings for LLM Training.